Part III: Answer the following questions as per the instructions given.

1.Given:

 P= Br. 10      and       TC = 120 + 4Q²

A. Find the profit maximizing level of price and quantity.

B. What will be the total profit?

2.Briefly explain the types of returns to scale.

3.Define market and State some of its elements.



1
Expert's answer
2022-02-08T10:36:34-0500

1)(a) Profit is maximum when: "MC=MR"

"TC=120+4Q^{2}"

"TR=P\times{Q}=10\times{Q}=10Q"

"MC=\frac{∆TC}{∆Q}=8Q"

"MR=\frac{∆TR}{∆Q}=10"

Equating MC and MR

"8Q=10"

"Q=1.25"

"Price=10Br"

(b) "Total Profit=TR-TC"

"(10Q)-(120+4Q^{2})"

"(10\u00d71.25)-(120+4\u00d71.25)"

"Profit\/Loss= -112.5"

2) Constant returns to scale is where an increase in inputs leads to a proportional increase in output.

Increasing returns to scale is where the output increases in a greater proportion than the increase in inputs.

Decreasing returns to scale is when output increases by a smaller proportion than the proportion increase in inputs.

3) A market is where buyers and sellers meet for exchange of goods and services. It is not necessarily a place.

Elements of a market include; product, price, place and promotion.